How do I solve
3x+2y=14
x-3y=-10?

Answer by subudear(62)

3x+2y=14 ......(1)
x-3y=-10 .......(2)
multiply equation (2) by 3 and subtract from equation (1)
3x+2y=14
3x-9y=-30
- + +
we get 11y = 44
y = 4
replace y = 4 in any eqaution, lets say (1) we get
3x + 2*4 = 14
x = 2
so solution is x = 2 and y =4